# Hard X-ray/Soft gamma-ray Laue Lenses for High Energy Astrophysics

### v1

@@
Abstract
The study of the celestial phenomena in the hard X-ray/soft gamma-ray band(20 keV–1 MeV) is very intriguing but also very difficult to be performed with the needed sensitivity. In this review I will discuss the astrophysical importance of the soft gamma-ray astronomy, its difficulties to solve its issues with the current instrumentation, and a possible solution achievable using focusing Laue lens. Concerning these instruments, I will discuss their functioning principle, how to achieve a high reflection efficiency, their imaging properties, the current feasibility studies, the technological developments and observation prospects.
@@ 1 Introduction
The only solution to the limitations of the current generation of gamma–ray instruments is the use of focusing instruments. As we will see, Laue lenses, based on diffraction from crystals in transmission configuration, can do that in the hard X–/soft gamma–ray ($<$ 1 MeV) domain and thus can provide a big leap in flux sensitivity with respect to the best non-focusing instruments. In addition, they can significantly improve the angular resolution with respect to mask telescopes (e.g., INTEGRAL IBIS).
Figure 1: The 158 keV line due to the radioactive 56Ni decay of the type Ia SN2014J, as observed with 150.24 ks INTEGRAL SPI observations, 3 weeks after the explosion. The line intensity is given in $10^{-4}$ photons cm-2s-1  units. Reprinted from Diehl14 .
@@ 2 Laue lens concept
[…]
@@ 2.2 Geometry of a Laue lens
A Laue lens is made of a large number of crystal tiles with diffracting planes in transmission configuration (see Fig. 2), disposed on a spherical cup of radius $R$, with their diffracting planes perpendicular to the sphere (spherical geometry, see Fig. 3). The focal spot is in the point 00 of the symmetry axis, at a distance $f=R/2$ from the cup, with $f$ being called focal length (see below). Eq. 2 requires that each crystal has to be oriented in such a way that the angle between the incident beam and the crystalline planes is given by the Bragg angle $\theta_{B}$. Thus, in order to focus a photon with energy $E$, for the first order diffraction which is the most efficient, the crystal has to be at a distance $r$ from the symmetry axis given by
@@ 2.3 Lens energy bandpass, inner and outer radii and role of the focal length
[…]
@@ 2.4 Crystal tile distribution in the lenses and their focusing
Given the spherical geometry of the lens, assuming that the crystal material and the diffracting planes ($hkl$) are the same for the entire lens, the best distribution of the crystal tiles on the spherical cup, for a high focal length ($>10$ m), is on concentric rings (see Fig. 4). Thus, for a given $r$, the diffracted energy $E$ is given by Eq. 4 for all crystals in each ring. For Laue lenses with small focal length ($<$ 10 m), in order to optimize the smoothness of the lens effective area versus energy (see below), the best geometrical distribution of the crystals on the spherical cup would be an Archimedes spiral (see, e.g., Pellicciotta06 ).
@@
Figure 4: Distribution, in concentric rings, of the crystals in a Laue lens with high focal length ($>10$ m). If the crystal tiles, for a ring of radius $r$, have the same diffracting planes ($hkl$), all of them reflect in the focus photons of energy $E$ given by Eq. 4. Left: focusing of mosaic crystals. Figure reprinted from vonBallmoos05 . Right: focusing of bent crystals.
@@ 3 Mosaic crystals and their reflectivity
$f(A)=\frac{B_{0}(2A)+|\cos 2\theta_{B}|\,B_{0}(2A|\cos 2\theta_{B}|)}{2A(1+\cos^{2}\theta_{B})}\$ (12)
@@ 3.1 Bent crystals and their reflectivity
As mentioned above, bent crystals with curvature radius equal to that of the lens surface, have the advantage of giving the best focusing of the reflected photons and a very smooth dependence of the effective area with energy.
In addition to these properties, the primary curvature produced from the crystal bending, can generate a curvature in the diffracting planes perpendicular to the crystal cross section. For particular crystallographic planes; in the other cases the diffracting planes, this effect is particularly important for the diffraction efficiency (ratio between the diffracted and the transmitted photons) and thus for the crystal reflectivity.
@@
Figure 5: X–ray diffraction in Laue geometry for flat (a) and bent (b) diffracting planes. Reprinted from Bellucci13
@@
Figure 6: Primary $\Omega_{p}$ and secondary or quasi-mosaic $\Omega_{s}$ bending angles in the case of (111) reflecting planes. The corresponding curvature radii are given by $R_{p}$ (lens curvature radius) and $R_{s}$ (secondary curvature radius). Reprinted from Guidi11 .
@@ 3.2 Techniques adopted for obtaining CDP crystals
CDP crystals can be obtained in various ways. One technique is the growing of a two-component crystal whose composition varies along the crystal growth axis, e.g., Si1-xGex Keitel99 ; Abrosimov05b . But the best techniques for Laue lenses, that provide also the best focusing, include the elastic bending of a perfect crystal tile (technique commonly adopted in synchrotron radiation facilities), the grooving of one of the surfaces of a crystal (working for thin ($\leq 2$ mm) crystals) Bellucci11 , the lapping of thin ($<$2 mm) crystals Ferrari13 , the deposition of a thin layer of carbon fiber onto mono-crystals Camattari14 .
Figure 8: Left panel: Diffracted and transmitted intensities as a function of the rocking angle for reflection (111) using synchrotron radiation at different energies. The transmitted intensity $I_{0}^{(0)}$ is normalized to 1. Reprinted from Keitel99 . Right panel: Measured 150 keV reflectivity of a Si(111) tile 2 mm thick that was bent by performing crossed grooves in one of its surfaces. The test was performed at the European Synchrotron Radiation Facility (ESRF) in Grenoble (France). Reprinted from Guidi13 .
@@ 3.3 Tests of quasi-mosaicity crystals
Several experimental tests of the quasi-mosaicity production have performed. In Fig. 8 (left panel) it is shown the measured reflectivity in the case of an ingot of Si1-xGex crystal, with $0.070.01<x<$ 0.07, grown along the $[111]$ direction with the Czockralski technique Keitel99 , while in the same figure (right panel) it is shown the reflectivity of Si(111) crystal tile, which was bent by grooving one of the cross sections Guidi13 . In both cases we can see very high reflectivities ($>$ 50%) and thus the formation of a QM structure.
@@ 4 Best crystal materials of a Laue lens
Figure 9: Density $N$ of a crystal unit cell versus atomic number Z of the elements. Reprinted from Ref. Frontera10 .
@@ 4.1 Best crystal thickness
In general, that implies a thickness of the order of 1 $\mu$m, which is not easy to get. For example, from extended tests performed on Cu(111) supplied by ILL Courtois05 , it results that the condition above is satisfied in single points Pellicciotta06 , but not when the entire crystal cross section is irradiated (values even higher than 100 $\mu$m have been found Barriere09 ).
@@ 5.1 Lens with mosaic crystals
In the case of mosaic crystals, the derived Point Spread Function (PSF) depends on the crystal size, on their mosaicity and on the accuracy of their positioning in the lens. Studies of the PSF with mosaic crystals have been performed by different authors (see, e.g., Pisa05b ; Barriere09c . In Fig. 11 (top panel) it is shown the on–axis PSF for a ring–shaped lens of 40 m focal length, 150–600 keV energy passband and a crystal tile cross section of 10$\times$10 mm2 Frontera08a . In this case it is assumed that the crystals have 1 arcmin mosaic spread and that they are properly oriented in the lens.
@@ 5.2 Lens with bent quasi-mosaic crystals
As discussed in section 2.4, the best focusing is obtained in the case of bent crystals with curvature radius equal to that of the spherical lens. In Fig. 12, top panel, it is shown the on-axis PSF of a Laue lens made of bent crystals of Si(111) and Ge(111) with quasi-mosaic diffracting planes, a passband from 50 to 600 keV and 20 m focal length Frontera21 , proposed for the ASTENA mission concept we will discuss later (see section 11). The expected angular resolution is 30 arcsec for a positioning accuracy of the crystal tiles $<$10 arcsec.
@@ 8.3 Laue lenses with tunable focal distance
The first tunable $\gamma$-ray lens prototype was developed and demonstrated by a collaboration project between Argonne National Laboratory (ANL) and the French space agency CNES Kohnle98 . The goal was that of observing more than one astrophysical line with a narrow band Laue lens. This goal requires the tuning of the Bragg angle and focal distance. The strategy adopted to change the Bragg angle was that of using piezo-driven actuators and an eddy-current sensor to determine the current position. In this way, an accuracy of 0.1--0.4 arcsec could be achieved. The drawback of this strategy is the large volume required, the decreased filling factor of crystal tiles, and the additional thickness that the reflected photons had to cross, in addition to the required power and the additional mass of the lens.
@@ 8.4 HAXTEL broad band Laue lens
The science goal of the HAXTEL (HArd X-ray TELescope) was that of building a broad band (70/100-600 keV) Laue lens. The crystals adopted were mosaic crystals of Cu(111) developed and produced at ILL Courtois05 . The tile cross-section was 15$\times$15 mm2 while its thickness was 2 mm. The mosaic spread of the crystals ranged from $\sim 2.5$ and $\sim 3.5$ arcmin Frontera08a .
@@
It was found that the PSF radius at which the expected fraction of focused photons reach the saturation (16 mm) corresponds to $\sim$60of the measured fraction. A better crystal orientation was later undertaken as discussed in Ferrari09 .
@@ 10 More recent projects of Laue lenses
Figure 14: Misalignment distribution of 48 crystal tiles used to build a Laue lens prototype, developed at University of California Berkeley. Reprinted from Wade15 .
@@ 10.1 Laue lens development at UC Berkeley
Goal of the project was to build a demonstrator of Laue lens, reflecting in the energy band from $\sim 90$ to $\sim 130$ keV, with mosaic crystals of Al(111), Fe(110) and Fe(200) Wade15 . For that, a dedicated X–ray beamline was developed (see details in Barriere14b and references therein). With the X–ray beam line the crystals were correctly oriented and mounted on the substrate. However, in spite that the positioning of the crystals within a few arcsec, given that the crystal diffracting planes were misaligned (±$\pm 1^{o}$) with respect to the cross sections, a wedge of glue was needed and thus a misalignment, due to shrinkage of the glue while curing, was found. The reported crystal misalignment distribution of the developed prototype is shown in Fig. 14.
@@ 10.2 The LAUE project
Figure 16: Misalignment distribution in Bragg angle of 11 bent crystal tiles (40 m curvature radius) of Ge(220) used to build a sector of Laue lens at the University of Ferrara. Adapted from Ferro24 .
@@ 10.3 Laue lenses based on Silicon Pore Optics technology
The Silicon crystal tiles are assumed to be been obtained by dicing commercial double-sided polished high purity Si wafers. Each of the tiles, with diffracting planes (111), is designed to be bent in azimuthal direction in order to fit the ring radius of the lens, while the Bragg direction (along the incident photon direction) is adjusted by inserting a proper wedge between contiguous crystal tiles in addition to an achievable radial bending of the crystal tiles. The crystal bending is expect to give a high diffraction efficiency. The ray tracing of a lens based on SiLCs was performed Girou17 with very good results in terms of focusing area ($<1$ mm2), while the study of such a Laue lens concept with a 20–158 keV passband and a focal of 8 m has been recently reported Barriere23 for polarimetry observations of celestial sources from a stratospheric balloon.
@@ 11 Prospects for Laue lens telescopes
As already mentioned (see, e.g., section 5), in 2019 a Laue lens telescope, designed for satisfying the above prescriptions, was proposed as one of the two instruments (NFT $=$ Narrow Field Telescope), the other being a broad-band Wide Field Monitor Imager and Spectrometer (WFM-IS), for the ASTENA mission concept proposed in two white papers to the European Space Agency (ESA) for its new long term programme ”Voyage 2050” Frontera21 ; Guidorzi21 . The lens telescope is made of bent crystals of Si(111) and Ge(111) that cover the nominal energy band from 50 to 600 keV with a still significant effective area up to 700 keV. A description of the lens and its performance, in addition to some key science goals that can be faced with NFT, can be found in Frontera21 . Other science goals that could be faced with ASTENA can be found in Guidorzi21 . The expected angular resolution and FOV of NFT have been given in section 5.2. Assuming the position sensitive detector described in section 6, the NFT expected sensitivity to continuum emission, to emission lines and to detect polarization of celestial soft gamma–ray sources is shown in Fig. 18. As far as the sensitivity to continuum emission, it is also shown the comparison of NFT with other experiments and missions. Among them, it is also included the broad-band (0.15 MeV–3 GeV) gamma–ray mission concept e-Astrogam Deangelis18 , that, in the common energy range with ASTENA NFT, makes use of a Compton telescope Deangelis18 . As can be seen from the top panel of Fig. 18, in this range ASTENA NFT is about two orders of magnitude more sensitive, but with a narrow FOV (4 arcmin). Thus NFT can perform deep studies of single sources and solve issues in which a high angular resolution is required, while e-Astrogam is suitable for monitoring broad sky fields, achieving a much better sensitivity than the current gamma–ray instruments by integrating on year times the source observations.

### v2

@@
11institutetext: Filippo Frontera 22institutetext: University of Ferrara, Department of Physics and Earth Sciences, Via Saragat, 1 - 44122 Ferrara, and INAF OAS, Via Gobetti 101, 40129 Bologna; 22email: filippo.frontera@unife.it
Hard X-ray/Soft gamma-ray Laue Lenses for High Energy Astrophysics
Filippo Frontera
@@ 1 Introduction
The only solution to the limitations of the current generation of gamma–ray instruments is the use of focusing instruments. As we will see, Laue lenses, based on diffraction from crystals in transmission configuration, can do that in the hard X–/soft gamma–ray (<< 1 MeV) domain and thus can provide a big leap in flux sensitivity with respect to the best non-focusing instruments. In addition, they can significantly improve the angular resolution with respect to mask telescopes (e.g., INTEGRAL IBIS).
Figure 1: The 158 keV line due to the radioactive 56Ni decay of the type Ia SN2014J, as observed with 150.24 ks INTEGRAL SPI observations, 3 weeks after the explosion. The line intensity is given in $10^{-4}$ photons cm-2s-1 units. Reprinted from Diehl14 . Reproduced with permission.
@@ 2 Laue lens concept
$\begin{aligned} d_{hkl}&=\frac{a}{\sqrt{h^{2}+k^{2}+l^{2}}}\qquad V=a^{3}\end{aligned}$ (1)
@@ 2.2 Geometry of a Laue lens
A Laue lens is made of a large number of crystal tiles with diffracting planes in transmission configuration (see Fig. 2), disposed on a spherical cup of radius $R$, with their diffracting planes perpendicular to the sphere (spherical geometry, see Fig. 3). The focal spot is in the point 0 of the symmetry axis, at a distance $f=R/2$ from the cup, with $f$ being called focal length (see below). Eq. 2 requires that each crystal has to be oriented in such a way that the angle between the incident beam and the crystalline planes is given by the Bragg angle $\theta_{B}$. Thus, in order to focus a photon with energy $E$, for the first order diffraction which is the most efficient, the crystal has to be at a distance $r$ from the symmetry axis given by
@@ 2.3 Lens energy bandpass, inner and outer radii and role of the focal length
$\begin{aligned} r_{min}\approx\frac{hcf}{d_{hkl}~E_{max}}\qquad r_{max}\approx\frac{hcf}{d_{hkl}~E_{min}}.\end{aligned}$ (5)
@@ 2.4 Crystal tile distribution in the lenses and their focusing
Given the spherical geometry of the lens, assuming that the crystal material and the diffracting planes (hklhkl) are the same for the entire lens, the best distribution of the crystal tiles on the spherical cup, for a high focal length (>10>10 m), is on concentric rings (see Fig. 4). Thus, for a given $r$, the diffracted energy $E$ is given by Eq. 4 for all crystals in each ring. For Laue lenses with small focal length (<< 10 m), in order to optimize the smoothness of the lens effective area versus energy (see below), the best geometrical distribution of the crystals on the spherical cup would be an Archimedes spiral (see, e.g., Pellicciotta06 ).
@@
Figure 4: Distribution, in concentric rings, of the crystals in a Laue lens with high focal length (>10>10 m). If the crystal tiles, for a ring of radius $r$, have the same diffracting planes (hklhkl), all of them reflect in the focus photons of energy $E$ given by Eq. 4. Left: focusing of mosaic crystals. Figure reprinted from vonBallmoos05 . Reproduced with permission. Right: focusing of bent crystals.
@@ 3 Mosaic crystals and their reflectivity
$f(A)=\frac{B_{0}(2A)+|\cos 2\theta_{B}|\,B_{0}(2A|\cos 2\theta_{B}|)}{2A(1+\cos^{2}\theta_{B})}\\$ (12)
@@ 3.1 Bent crystals and their reflectivity
As mentioned above, bent crystals with curvature radius equal to that of the lens surface have the advantage of giving the best focusing of the reflected photons and a very smooth dependence of the effective area with energy.
In addition to these properties, the primary curvature produced from the crystal bending can generate a curvature in the diffracting planes perpendicular to the main crystal cross section for particular crystallographic planes. This effect is particularly important for the diffraction efficiency (ratio between the diffracted and the transmitted photons) and thus for the crystal reflectivity.
@@
Figure 5: X–ray diffraction in Laue geometry for flat (a) and bent (b) diffracting planes. Reprinted from Bellucci13 . Reproduced with permission.
@@
Figure 6: Primary $\Omega_{p}$ and secondary or quasi-mosaic $\Omega_{s}$ bending angles in the case of (111) reflecting planes. The corresponding curvature radii are given by $R_{p}$ (lens curvature radius) and $R_{s}$ (secondary curvature radius). Reprinted from Guidi11 . Reproduced with permission.
@@ 3.2 Techniques adopted for obtaining CDP crystals
CDP crystals can be obtained in various ways. One technique is the growing of a two-component crystal whose composition varies along the crystal growth axis, e.g., Si1-xGex Keitel99 ; Abrosimov05b . But the best techniques for Laue lenses, that provide also the best focusing, include the elastic bending of a perfect crystal tile (technique commonly adopted in synchrotron radiation facilities), the grooving of one of the surfaces of a crystal (working for thin ($\leq 2$ mm) crystals) Bellucci11 , the lapping of thin (<<2 mm) crystals Ferrari13 , the deposition of a thin layer of carbon fiber onto mono-crystals Camattari14 .
Figure 8: Left panel: Diffracted and transmitted intensities as a function of the rocking angle for reflection (111) using synchrotron radiation at different energies. The transmitted intensity $I_{0}^{(0)}$ is normalized to 1. Reprinted from Keitel99 . Reproduced with permission. Right panel: Measured 150 keV reflectivity of a Si(111) tile 2 mm thick that was bent by performing crossed grooves in one of its surfaces. The test was performed at the European Synchrotron Radiation Facility (ESRF) in Grenoble (France). Reprinted from Guidi13 . Reproduced with permission.
@@ 3.3 Tests of quasi-mosaicity crystals
Several experimental tests of the quasi-mosaicity production have performed. In Fig. 8 (left panel) it is shown the measured reflectivity in the case of an ingot of Si1-xGex crystal, with $0.01<x<0.07$, grown along the $[111]$ direction with the Czockralski technique Keitel99 , while in the same figure (right panel) it is shown the reflectivity of Si(111) crystal tile, which was bent by grooving one of the cross sections Guidi13 . In both cases we can see very high reflectivities (>> 50%) and thus the formation of a QM structure.
@@ 4 Best crystal materials of a Laue lens
Figure 9: Density NN of a crystal unit cell versus atomic number Z of the elements. Reprinted from Ref. Frontera10 .
@@ 4.1 Best crystal thickness
[…]
@@ 5.1 Lens with mosaic crystals
In the case of mosaic crystals, the derived Point Spread Function (PSF) depends on the crystal size, on their mosaicity and on the accuracy of their positioning in the lens. Studies of the PSF with mosaic crystals have been performed by different authors (see, e.g., Pisa05b ; Barriere09c . In Fig. 11 (top panel) it is shown the on–axis PSF for a ring–shaped lens of 40 m focal length, 150–600 keV energy passband and a crystal tile cross section of $\times10$ mm2 Frontera08a . In this case it is assumed that the crystals have 1 arcmin mosaic spread and that they are properly oriented in the lens.
@@ 5.2 Lens with bent quasi-mosaic crystals
As discussed in section 2.4, the best focusing is obtained in the case of bent crystals with curvature radius equal to that of the spherical lens. In Fig. 12, top panel, it is shown the on-axis PSF of a Laue lens made of bent crystals of Si(111) and Ge(111) with quasi-mosaic diffracting planes, a passband from 50 to 600 keV and 20 m focal length Frontera21 , proposed for the ASTENA mission concept we will discuss later (see section 11). The expected angular resolution is 30 arcsec for a positioning accuracy of the crystal tiles <<10 arcsec.
@@ 8.3 Laue lenses with tunable focal distance
The first tunable $\gamma$-ray lens prototype was developed and demonstrated by a collaboration project between Argonne National Laboratory (ANL) and the French space agency CNES Kohnle98 . The goal was that of observing more than one astrophysical line with a narrow band Laue lens. This goal requires the tuning of the Bragg angle and focal distance. The strategy adopted to change the Bragg angle was that of using piezo-driven actuators and an eddy-current sensor to determine the current position. In this way, an accuracy of 0.1-0.4 arcsec could be achieved. The drawback of this strategy is the large volume required, the decreased filling factor of crystal tiles, and the additional thickness that the reflected photons had to cross, in addition to the required power and the additional mass of the lens.
@@ 8.4 HAXTEL broad band Laue lens
The science goal of the HAXTEL (HArd X-ray TELescope) was that of building a broad band (70/100-600 keV) Laue lens. The crystals adopted were mosaic crystals of Cu(111) developed and produced at ILL Courtois05 . The tile cross-section was $\times15$ mm2 while its thickness was 2 mm. The mosaic spread of the crystals ranged from $\sim 2.5$ and $\sim 3.5$ arcmin Frontera08a .
@@
It was found that the PSF radius at which the expected fraction of focused photons reach the saturation (16 mm) corresponds to $\sim$60% of the measured fraction. A better crystal orientation was later undertaken as discussed in Ferrari09 .
@@ 10 More recent projects of Laue lenses
Figure 14: Misalignment distribution of 48 crystal tiles used to build a Laue lens prototype, developed at University of California Berkeley. Reprinted from Wade15 . Reproduced with permission.
@@ 10.1 Laue lens development at UC Berkeley
Goal of the project was to build a demonstrator of Laue lens, reflecting in the energy band from $\sim 90$ to $\sim 130$ keV, with mosaic crystals of Al(111), Fe(110) and Fe(200) Wade15 . For that, a dedicated X–ray beamline was developed (see details in Barriere14b and references therein). With the X–ray beam line the crystals were correctly oriented and mounted on the substrate. However, in spite that the positioning of the crystals within a few arcsec, given that the crystal diffracting planes were misaligned ($\pm 1^{o}$) with respect to the cross sections, a wedge of glue was needed and thus a misalignment, due to shrinkage of the glue while curing, was found. The reported crystal misalignment distribution of the developed prototype is shown in Fig. 14.
@@ 10.2 The LAUE project
Figure 16: Misalignment distribution in Bragg angle of $1$ bent crystal tiles (40 m curvature radius) of Ge(220) used to build a sector of Laue lens at the University of Ferrara. Adapted from Ferro24 .
@@ 10.3 Laue lenses based on Silicon Pore Optics technology
The Silicon crystal tiles are assumed to be been obtained by dicing commercial double-sided polished high purity Si wafers. Each of the tiles, with diffracting planes (111), is designed to be bent in azimuthal direction in order to fit the ring radius of the lens, while the Bragg direction (along the incident photon direction) is adjusted by inserting a proper wedge between contiguous crystal tiles in addition to an achievable radial bending of the crystal tiles. The crystal bending is expect to give a high diffraction efficiency. The ray tracing of a lens based on SiLCs was performed Girou17 with very good results in terms of focusing area (<1<1 mm2), while the study of such a Laue lens concept with a 20–158 keV passband and a focal of 8 m has been recently reported Barriere23 for polarimetry observations of celestial sources from a stratospheric balloon.
@@ 11 Prospects for Laue lens telescopes
As already mentioned (see, e.g., section 5), in 2019 a Laue lens telescope, designed for satisfying the above prescriptions, was proposed as one of the two instruments (NFT == Narrow Field Telescope), the other being a broad-band Wide Field Monitor Imager and Spectrometer (WFM-IS), for the ASTENA mission concept proposed in two white papers to the European Space Agency (ESA) for its new long term programme ”Voyage 2050” Frontera21 ; Guidorzi21 . The lens telescope is made of bent crystals of Si(111) and Ge(111) that cover the nominal energy band from 50 to 600 keV with a still significant effective area up to 700 keV. A description of the lens and its performance, in addition to some key science goals that can be faced with NFT, can be found in Frontera21 . Other science goals that could be faced with ASTENA can be found in Guidorzi21 . The expected angular resolution and FOV of NFT have been given in section 5.2. Assuming the position sensitive detector described in section 6, the NFT expected sensitivity to continuum emission, to emission lines and to detect polarization of celestial soft gamma–ray sources is shown in Fig. 18. As far as the sensitivity to continuum emission, it is also shown the comparison of NFT with other experiments and missions. Among them, it is also included the broad-band (0.15 MeV–3 GeV) gamma–ray mission concept e-Astrogam Deangelis18 , that, in the common energy range with ASTENA NFT, makes use of a Compton telescope Deangelis18 . As can be seen from the top panel of Fig. 18, in this range ASTENA NFT is about two orders of magnitude more sensitive, but with a narrow FOV (4 arcmin). Thus NFT can perform deep studies of single sources and solve issues in which a high angular resolution is required, while e-Astrogam is suitable for monitoring broad sky fields, achieving a much better sensitivity than the current gamma–ray instruments by integrating on year times the source observations.